OpenStudy (anonymous):

How can 2x2 = x2 + 4x + 8 be set up as a system of equations? y = x2 - 4 y = x2 + 2x + 8 y = x2 y = x2 + 4x + 8 y = 2x2 y = x2 + 4x + 8 y = x2 y = 2x + 8

OpenStudy (cggurumanjunath):

y = 2x2 y = x2 + 4x + 8

OpenStudy (anonymous):

Thanks! @CGGURUMANJUNATH I have another question exactly like this that I don't know either.

OpenStudy (anonymous):

How can x + 3 = -2x + 5 be set up as a system of equations? y = x + 13 y = -2x + 5 y = x + 3 y = -3x + 11 y = x + 12 y = -2x + 1 y = x + 3 y = -2x + 5

OpenStudy (cggurumanjunath):

can u guess similarly now @tornjeansxo ?

OpenStudy (anonymous):

Is it B or D?

OpenStudy (cggurumanjunath):

u tell me !

OpenStudy (anonymous):

I think it's D, haha

OpenStudy (cggurumanjunath):

:)

OpenStudy (anonymous):

thanks so much!

OpenStudy (cggurumanjunath):

wc.